OpenStudy (anonymous):

The graph below represents which system of inequalities? graph of two infinite lines that intersect at a point. One line is solid and goes through the points negative 3, 0, negative 4, negative 1 and is shaded in below the line. The other line is solid, and goes through the points 1, 1, 2, negative 1 and is shaded in below the line. y ≤ −2x + 3 y ≤ x + 3 y ≥ −2x + 3 y ≥ x + 3 y ≤ −3x + 2 y ≤ −x + 2 y > −2x + 3 y > x + 3

OpenStudy (anonymous):

can you help me

OpenStudy (phi):

can you find the slope between points negative 3, 0, negative 4, negative 1 which I assume means (-3,0) and (-4, -1)

OpenStudy (phi):

find "change in y" divided by "change in x"

OpenStudy (anonymous):

so it would look like -1-0/-4+3

OpenStudy (phi):

yes. now simplify

OpenStudy (anonymous):

its 1

OpenStudy (phi):

if you write the equation of a line in the form y = mx + b m is the slope in this case y = 1x + b or just y= x+ b when they say One line is solid that means either \( \le \) or \( \ge\)

OpenStudy (anonymous):

ok so the answer is D

OpenStudy (phi):

is shaded in below the line that means the "y" values are less than or equal to the line so you should look for \[ y \le x + b \] as one of the answers. cross off any choice that does not have that as an option.

OpenStudy (anonymous):

is the B three

OpenStudy (phi):

choice D y > −2x + 3 y > x + 3 means: 1) the > means the line is dotted. 2) the shaded area is y bigger (i.e. above) the line so D is wrong two different ways.

OpenStudy (anonymous):

ok so its A

OpenStudy (phi):

we could figure out what b is in y = x + b by using one of the points: (-3,0) replace x with -3 and y with 0 0 = -3 + b add +3 to both sides +3 = -3 + 3 + b or 3= b so yes , b is 3

OpenStudy (phi):

Yes. the answer is A B has the wrong >= sign and C does not have y <= x + 3

OpenStudy (phi):

If you wanted, you could find the slope of the second line, and figure out its equation. But unless there is a typo, it will be y <= -2x+3

OpenStudy (anonymous):

thanks for your help and have a great fourth of july